OpenStudy (anonymous):

what two numbers you can multiply and get 64 and add or subtract to get 24?

OpenStudy (anonymous):

AB=64, A+B=24; A=24-B, (24-B)B=64, 24B-B^2-64=0, quadratic formula. You can do it from here, right?